Help solve:
average tuition in 1990 was $2035.00. By 2000 it had risen approx 85%. What was the approx. cost in 2000

Found 2 solutions by consc198, ikleyn:
Answer by consc198(59)

85 divided by 100 = 0.85
0.85 x $2035 = $1729.75
Approx. Cost = $1700

Answer by ikleyn(53937)

.
Help solve:
average tuition in 1990 was $2035.00. By 2000 it had risen approx 85%. What was the approx. cost in 2000
~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution by @consc198 in his post is, OBVIOUSLY, wrong.
        For a correct solution see what follows.


The cost in 2000, according to the problem is


    2035 + 0.85*2035 = 1.85*2035 = 3764.75 dollars.   


Since they ask "approx.", you may round it to 3765 dollars.

Solved correctly.

Do not trust to this person, @consc198, since he regularly produces wrong solutions without any hesitation.


Simply saying, a person, whose level in Math is about 3 of 5, decided that he can and he should teach others.